### v1

# Look Before you Leap: Estimating LLM Benchmark Scores from Descriptions

Abstract
Progress in large language models is constrained by an evaluation bottleneck: build a benchmark, evaluate models and settings, then iterate. We therefore ask a simple question: can we forecast outcomes before running any experiments? We study text-only performance forecasting: estimating a model’s score from a redacted task description and intended configuration, with no access to dataset instances. To support systematic study, we curate Precog, a corpus of redacted description–performance pairs spanning diverse tasks, domains, and metrics. Experiments show the task is challenging but feasible: models equipped with a retrieval module that excludes source papers achieve moderate prediction performance with well-calibrated uncertainty, reaching mean absolute error as low as $8.7$ on the Accuracy subset at high-confidence thresholds. Our analysis indicates that stronger reasoning models engage in diverse, iterative querying, whereas current open-source models lag and often skip retrieval or gather evidence with limited diversity. We further test a zero-leakage setting—forecasting on newly released datasets or experiments before their papers are indexed—where GPT-5 with built-in web search still attains non-trivial prediction accuracy. Overall, our corpus and analyses offer an initial step toward open-ended anticipatory evaluation, supporting difficulty estimation and smarter experiment prioritization. [^1] Our data and code and are available in https://github.com/JJumSSu/PRECOG
1 Introduction
Modern LLM development is evaluation-centric: define a benchmark, sweep models and configurations, iterate (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Team et al., 2023). As benchmarks proliferate, exhaustive curation and full sweeps demand substantial compute, researcher time, and funding (Mazumder et al., 2023). We therefore seek an ex ante performance forecaster to help designers decide what to build and where to measure. For example, consider a team proposing a binary “solver-choice” task on an ordinary differential equations benchmark (Figure 1; Gaonkar et al., 2025). Before annotating a single example, they can ask a concrete what-if: will GPT-4.1 (Achiam et al., 2023) without CoT (Wei et al., 2022) land nearer 30 or 60 Accuracy points? A forecast like this has the potential to quickly inform scope and spend—how large a pilot to run, which baselines to include, and whether the task merits full-scale annotation.
Figure 1: Precog pairs a redacted task/dataset description with its target metric. Given the inputs (a redacted description of dataset and an experimental configuration (Gaonkar et al., 2025)), the goal is to predict the target metric. The regressor conditions on performance-relevant details and may retrieve corroborating evidence from the literature.
Two adjacent lines of work on ML model performance prediction address related goals: (i) missing-entry prediction, which infers unobserved cells from observed ones (Xia et al., 2020; Ye et al., 2021; 2023; Zhao et al., 2025); and (ii) efficient evaluation, which extrapolates full-benchmark performance from a small set of pilot runs (Maia Polo et al., 2024; Pacchiardi et al., 2024; Anugraha et al., 2025). Both assume seed measurements or closely related tasks and typically rely on engineered numeric features. In contrast, we target a zero-example, open-ended setting in which retrieval is driven by natural-language task descriptions, supporting researchers who are creating new datasets or designing new experiments. Prior work has also explored the capability of LLMs to forecast the outcomes of future geopolitical events in online prediction markets (Karger et al., 2024; Halawi et al., 2024), however to the best of our knowledge, this is the first work to explore the capability of LLMs to predict their own performance on a new task based only on a text description.
We formulate text-only performance forecasting as regression: given a natural-language specification of the dataset and evaluation setup, including task, data provenance, prompting/decoding choices, and metric/protocol, we predict a target metric normalized to a 0100100 scale. As in Figure 1, the input is a description of the proposed task and configuration; the output is a predicted metric value (e.g., $32$). Unlike feature-engineered baselines, we operate directly on free-form, literature-style text and optionally use retrieval to consult related prior work, rather than interpolation from seed results.
To study this setting systematically, we curate Precog, a corpus of arXiv-mined experimental records (Park et al., 2025) that, for each record, pairs the dataset-source paper (introducing the benchmark) with the result-reporting paper (reporting the score). From these sources, we synthesize redacted descriptions that preserve task and setup details while masking dataset names and numeric results. At retrieval time, we exclude each record’s own source papers to limit leakage. To further probe residual leakage risks (Tian et al., 2024), we also evaluate a streaming regime that ingests arXiv releases and makes predictions before they are indexed by search engines.
We then evaluate zero-shot LLM regressors, optionally augmented with literature search. Given a redacted description, the model may iteratively issue evidence-seeking queries (ReAct-style (Yao et al., 2023)) and outputs a point estimate alongside a self-reported confidence (Yoon et al., 2025). We assess calibration by sweeping a confidence threshold and reporting the coverage–MAE trade-off, and we analyze search tool-use behavior.
On Precog, a proprietary reasoning model with literature retrieval attains MAE $14.0$ across seven metric groups and $8.7$ on the Accuracy subset at high-confidence thresholds, with reasonably calibrated uncertainty—indicating the task is hard but tractable. The proprietary reasoning model employs diverse, iterative querying, whereas current open-source models underperform and exhibit shallower tool use. In streaming prediction, GPT-5 (OpenAI, 2025) with built-in web search delivers competitive, though imperfect, performance on newly released datasets.
Altogether, our study shows that LLM-based regression models, which predict model performance from text descriptions, can help inform benchmark design, prioritize configurations, determine dataset sizes, and guide pilot experiment planning—while reducing the compute and time required for exploratory analysis. We release Precog, a redacted-description benchmark grounded in scientific literature, along with the accompanying prompts, protocols, and experimental code. We hope these resources and our evaluation methodology will support future advances in predictive model development and evaluation.
2 Related Work
Performance Prediction
Prior work predicts missing table entries from observed cells plus side information, using matrix-completion or response-surface models (Achille et al., 2019; Xia et al., 2020; Chen & Ritter, 2021; Ye et al., 2021; 2023; Rabinovich et al., 2023; Shi et al., 2025; Zhao et al., 2025). These methods typically assume (i) seed results on related tasks, (ii) stable task identities, and (iii) overlap that enables interpolation. A complementary line of work targets efficient evaluation: execute a small subset of measurements and extrapolate full-benchmark performance via learning-curve or proxy modeling (Maia Polo et al., 2024; Pacchiardi et al., 2024; Anugraha et al., 2025). These approaches still require running some evaluations on the target (or closely related) benchmark.
LLMs for Regression
Recent work explores LLMs as regressors in structured experimental settings, such as AutoML or black-box optimization (Song et al., 2024; Vacareanu et al., 2024; Nguyen et al., 2024; Requeima et al., 2024; Tang et al., 2024). OmniPred (Song et al., 2024), for instance, trains LLMs on millions of hyperparameter optimization trajectories from Google’s internal Vizier database and demonstrates strong regression accuracy across diverse tasks. However, these approaches rely on structured parameterizations (e.g., key–value hyperparameter strings) and assume access to seed trials or finetuning on the target task. In contrast, we study a zero-example setting where the sole input is a free-form natural-language description of a novel dataset and evaluation protocol, supporting researchers who have not yet collected any results.
Extracting Structured Records from Litertaure
Prior work extracts structured result tuples (e.g., task, dataset, metric) from scientific articles (Singh et al., 2019; Hou et al., 2019; Kardas et al., 2020; Yang et al., 2022; Bai et al., 2023; Şahinuç et al., 2024; Park et al., 2025). Precog builds on these tuples: for each experimental record, we synthesize a richer natural-language description that augments the tuple with additional context.
3 Task
Formulation
We study zero-example, text-only performance forecasting as a regression problem. Given a self-contained, redacted description $x_{i}$ of a dataset and evaluation configuration, the model performs reasoning and outputs a numeric estimate $\hat{y}_{i}\in[0,100]$. We treat $x_{i}$ as a structured yet free-form text description that captures contextual nuances of the task, dataset, and experimental configuration without requiring access to instances. To compare heterogeneous metrics (e.g., Accuracy, F1, BLEU), we convert each paper-reported score $y_{i}^{\text{paper}}$ to a common 0–100 scale via a simple normalization $(N:\mathbb{R}\!\to\![0,100])$; $y_{i}=N(y_{i}^{\text{paper}})$. Concretely, values already on a 0–100 scale remain unchanged, while fractional scores in $([0,1])$ are multiplied by 100 (e.g., Accuracy $(0.35\mapsto 35.0)$; F1 $(83.2\mapsto 83.2)$.
Given a redacted description $x_{i}$, an LLM regressor $f_{\theta}$ predicts the normalized target as $\hat{y}_{i}=f_{\theta}(x_{i},Z_{i})$, where $Z_{i}$ is an optional set of retrieved evidence. We obtain $Z_{i}$ by querying a time-bounded literature corpus $\mathcal{C}$ with a top-$k$ retriever $R_{k}$ on a query formed from $x_{i}$, i.e., $Z_{i}=R_{k}(q(x_{i}),\,\mathcal{C})$. Setting $k=0$ yields the description-only variant $f_{\theta}(x_{i},\varnothing)$. At inference, $f_{\theta}$ may iterate in a ReAct-style loop—reason (Yao et al., 2023), issue $q(\cdot)$, retrieve $Z_{i}$, refine—and then output $\hat{y}_{i}$. For example, for the open-domain QA description above, $f_{\theta}$ can retrieve prior exact-match (EM) ranges for similar setups and use them to refine $\hat{y}_{i}$.
Evaluation
Following prior work (Srinivasan et al., 2021; Ahuja et al., 2022; Jawahar et al., 2024), we evaluate $f_{\theta}$ with two standard criteria: mean absolute error (MAE) and the Pearson correlation $\mathrm{r}$ between $y_{i}$ and $\hat{y}_{i}$. MAE (lower is better) measures absolute accuracy across heterogeneous metrics, while $\mathrm{r}$ (higher is better) assesses whether predictions preserve the relative ordering of tasks/configurations. Used together, they capture both value accuracy and trend fidelity despite differences in metric scales and variances—an evaluation protocol that is standard for performance prediction.
Task Design
Work Text-Desc. Feat. Retrieval Runs
Xia et al. (2020)
Ye et al. (2021)
Ye et al. (2023)
Maia Polo et al. (2024)
Pacchiardi et al. (2024)
Anugraha et al. (2025)
Precog (ours)
Table 1: Comparison of performance-prediction schemas. Text-Desc.: free-form descriptions. Feat.: engineered numerics features. Retrieval: literature search at prediction time. Runs: target model run on a subset.
Difference from Prior Work
Table 1 contrasts our setting with prior work: (i) it is literature-grounded and open-ended, requiring retrieval rather than seed-set evaluation; and (ii) it predicts from free-form natural-language descriptions rather than engineered numeric features. Operationally, you can onboard new tasks without running target models or maintaining seed evaluations. Free-form descriptions, augmented by retrieval from external evidence sources, capture protocol nuances and handle open-ended variation, while yielding traceable evidence for human-in-the-loop review.
4 Data Curation
To evaluate at scale, we need a large number of diverse and realistic, open-ended pairs $(x_{i},y_{i})$. Rather than synthesizing experimental configurations and running automated experiments, which is both costly and unlikely to capture the kinds of problems practitioners actually investigate, we semi-automatically curate experimental configurations and their reported results directly from the scientific literature. Our pipeline harvests redacted task descriptions and reported scores from published papers; humans intervene only for precision-critical steps. We follow the pipeline in Figure 2: (i) collect arXiv experimental records (section 4.1); (ii) define the record provenance and paper types (section 4.2); (iii) recover each record’s sources (section 4.3); (iv) extract a schema-aligned, redacted description (section 4.4); and (v) perform validation and quality control (section 4.5).
4.1 Source Corpus
We build off of LLMEvalDB (Park et al., 2025), a structured arXiv-mined repository (2023–2024) dominated by non-finetunable, proprietary models—suited to isolating dataset/task effects with less model/method variance. Each experimental record is a table cell with the dataset name, configuration context, and reported metric. Values are normalized to define $y_{i}$. We consider seven metrics with natural [0,100] scales: Accuracy, F1, Recall, Precision, ROUGE, BLEU, and EM.
4.2 Record Provenance and Paper Types
Each record is grounded in two literature sources: the result paper, which reports the score, and the dataset paper, which introduced the dataset. These may coincide or differ. For example, Paper A introduces a QA dataset and reports accuracy for several models (result paper = dataset paper). Later, Paper B evaluates a new prompting/evaluation protocol on the same dataset under a revised setup (result paper $\neq$ dataset paper) (Park et al., 2021; Rangapur & Rangapur, 2024).
4.3 Dataset Source Paper Retrieval
To compile faithful descriptions for each record, we first collect its sources. LLMEvalDB (Park et al., 2025) provides the result papers. When the result paper differs from the dataset paper, we automatically recover the latter by inferring the citation with an LLM, resolving it to an arXiv ID, downloading the PDF, and verifying the dataset name/title. We then run a rule-based check to confirm that the PDF explicitly mentions the dataset (string match). Because source fidelity is critical, any remaining errors are corrected via targeted human edits.
4.4 Extracting Redacted Input Description
Given the sources for the result and dataset papers, an extractor LLM produces a schema-aligned, redacted description $x_{i}$ that summarizes the task (problem definition and input/label format), data collection (provenance, domain, collection protocol), evaluation (metric and protocol), difficulty cues (number of classes, long context, reasoning depth), the relevant subset/split, the prompting strategy (zero/few-shot, CoT), and any other record-specific settings (context-window, tool/retrieval use), when reported; see the Appendix D for illustrative examples.
We systematically remove the dataset names, performance numbers, and statistics. To further reduce extraction-time leakage, a second LLM rewriting pass is applied and the same checks are re-enforced, yielding faithful, self-contained, anonymized inputs $x_{i}$.
Figure 2: Overview of the Precog curation pipeline. For each record, we combine the result paper (Mitchell et al., 2023), which reports the score and specifies the evaluation protocol and configuration, with the dataset paper (Moskvichev et al., 2023), which provides the task description, to produce a schema-aligned input. The score $0.30$ reported is normalized to serve as the target.
4.5 Data Statistics and Quality Control
Statistic Value
Record counts
Total instances 767
result paper == dataset paper 504
result paper $\neq$ dataset paper 263
Unique counts
Unique datasets 528
result paper $\cup$ dataset paper 631
Human validation (30-sample audit)
Anonymous (0–1) 1.00
Schema coverage (1–5) 4.61
Source support (1–5) 4.87
Table 2: Precog statistics and human validation of extracted descriptions.
We construct a bounded yet diverse corpus via stratified sampling over source papers, enforcing caps per paper and per dataset plus per-source quotas to prevent any single source from dominating. This yields 767 instances spanning 528 distinct datasets and 631 unique papers, covering many datasets and configurations rather than a single benchmark (Ye et al., 2023). Rather than scaling the number of instances—which would be costly for reasoning models—we focused on curating diverse and high-quality examples using the GPT-5 (OpenAI, 2025) API (\$200) and 36 hours of human validation.  [^2] We also release $\sim 9\mathrm{k}$ pre-extraction intermediate sources to support scaling evaluation and training.
Because Precog is LLM-assisted, we add fidelity safeguards: (i) humans verify and correct each record’s dataset–source link; (ii) dataset names are canonicalized and we sample over unique source pairs to promote diversity; and (iii) metrics and related fields inherited from LLMEvalDB (Park et al., 2025) exhibit near-perfect extraction accuracy. We further audited 30 random descriptions under a fixed protocol—checking anonymization (binary: no dataset names or performance numbers), schema coverage (5-point), and source support vs. the PDFs (5-point). Results (Table 2) show well-anonymized, accurate, and comprehensively grounded descriptions, suitable as schema-aligned, leakage-resistant inputs for text-only performance forecasting.
5 Experiments
Baselines
We compare against three families of baselines. First, we report the overall test-split mean across all metrics (Section 5.1) and the per-metric test-split means (Section 5.2). Because no training or validation set is available, the test-set average serves as a strong baseline for comparison. Second, we evaluate text-encoder–based regressors. We embed each description and train classical models—k-nearest neighbors (Guo et al., 2003) and XGBoost (Chen & Guestrin, 2016)—on these embeddings. With no dedicated training set, we use k-fold cross-validation (Fushiki, 2011): fit on k-1 folds, evaluate on the held-out fold, and average over folds. For embeddings, we use E5-Mistral-7B (Wang et al., 2023), a leading open-source encoder.
LLMs
We evaluate both proprietary and open-source models in reasoning and non-reasoning settings to isolate the contribution of latent chain-of-thought style reasoning from raw generation quality. On the proprietary side we use GPT-5 (OpenAI, 2025) as a reasoning model. On the open-source side we include reasoning-enabled model Qwen3-32B (Yang et al., 2025), alongside non-reasoning models Qwen2.5-32B, Qwen2.5-72B (Team, 2024). Using matched prompts and decoding settings across these pairs lets us attribute gains specifically to reasoning capabilities. We prompt the model to output its final estimate in boxed{} with a brief rationale, then parse the boxed value during post-processing. Predictions are produced via greedy decoding with default generation settings.
Search Retrieval
All models share a single custom retriever over the arXiv API [^3] https://info.arxiv.org/help/api/index.html that returns canonical arXiv PDF links. To prevent leakage, we exclude each record’s source papers using identifier during retrieval. We initially tested GPT-5’s built-in web search—more token-efficient but corpus-uncontrolled—which frequently surfaced source papers; in an audit, $\sim\!70\%$ of evaluation dataset-source papers became identifiable, undermining realism. Consequently, despite higher processing cost, we rely on the custom retriever for GPT-5. We report results with retrieval on and off to assess gains from literature grounding beyond the model’s prior. Each query returns at most two documents, and the agent may issue up to four retrievals under ReAct-style iterative prompting (Yao et al., 2023), implemented via Qwen-Agent  [^4] https://github.com/QwenLM/Qwen-Agent and OpenAI Response API  [^5] https://platform.openai.com/docs/api-reference/responses.
5.1 Evaluation Results
Model MAE $\mathbf{r}$
Baselines
Test set average 21.4
E5-Mistral + kNN 19.9 0.30
E5-Mistral + XGBoost 20.3 0.30
LLMs
GPT-5 14.7 0.69
GPT-5 + Search 14.0 0.73
Qwen3-32B 19.6 0.40
Qwen3-32B + Search 20.8 0.31
Qwen2.5-32B 23.5 0.40
Qwen2.5-32B + Search 22.0 0.24
Qwen2.5-72B 22.4 0.44
Qwen2.5-72B + Search 21.4 0.34
Table 3: Aggregated results (767 instances); Search = arXiv retrieval. indicates significance at $p<0.05$ (one-sided binomial sign test).
We report overall results in Table 3. Except for GPT-5 (OpenAI, 2025), no baseline meaningfully outperforms the test-set global mean. GPT-5 with Search achieves the best overall MAE (14.0) yet still leaves substantial headroom, while GPT-5 without Search also performs well, suggesting that pretrained knowledge and difficulty cues already carry useful signal. Classical text-encoder baselines—kNN and XGBoost with k-fold CV—do not meaningfully outperform the global mean. This suggests that effective forecasting most likely requires (i) reasoning, (ii) representations that capture fine-grained cues in long descriptions, and (iii) knowledge beyond simple interpolation from seed results. Among Qwen variants (Team, 2024; Yang et al., 2025), Qwen3-32B (reasoning) leads, however enabling Search gives inconsistent MAE and can degrade it, indicating brittle tool use. Qwen2.5-72B outperforms Qwen2.5-32B but still trails Qwen3-32B (reasoning).
5.2 Further Analysis Across Data Subsets
We report per-metric MAE against a test-calibrated per-metric mean baseline—stronger than a single global mean—and compare GPT-5 (with/without Search) across seven metric groups, given its accurate, non-trivial forecasts. Figure 3 reports per-metric MAE. GPT-5 generally outperforms the metric-calibrated test-set mean baseline. The Accuracy subset—our largest slice—mirrors the overall trend. By contrast, for BLEU and ROUGE, whose gold scores have low variance, LLMs sometimes fall below the baseline, reflecting the difficulty of forecasting fine-grained metric values.
Peer-reviewed arXiv
# Samples 520 247
GPT-5 14.6 14.5
GPT-5 + Search 13.8 14.4
Table 4: Peer-reviewed vs. arXiv: sample counts and MAE.
Figure 3: Per-metric MAE (↓) by model, grouped by evaluation metric. The red bar marks the per-metric average with its absolute MAE label; other bars show delta to that average. The test set comprises 767 instances: 419 Accuracy, 152 F1, 56 Recall, 42 Precision, 35 EM, 32 ROUGE, and 31 BLEU.
We further stratify results by venue to assess whether publication status correlates with predictability. We compare two groups: Peer-reviewed samples appearing in refereed venues, arXiv postings in CoRR preprints. As shown in Table 4, the peer-reviewed subset and arXiv subset exhibit very similar MAE for GPT-5. Adding Search gives modest gains on the peer-reviewed subset vs. arXiv, hinting at answer leakage; accordingly, we also evaluate a streaming setting (Section 7) that forecasts on newly released datasets before citations/indexing. Overall differences are small, suggesting venue has limited effect and task characteristics dominate.
5.3 Human prediction results
Model MAE $\mathrm{r}$
Human 19.6 0.40
Test subset average 21.7
GPT-5 + Search 12.3 0.79
Table 5: Human Prediction Results on Accuracy subset ($n=140$).
In the previous sections, we demonstrated that LLMs can generate accurate predictions from dataset descriptions. A natural next question is how these predictions compare to those made by human experts. To explore this, we asked six graduate students and one professor of computer science, all specializing in NLP and machine learning, to each predict 20 accuracies based solely on the same text descriptions provided to models. On this 140-item subset, human predictions correlated moderately with ground truth ($r=0.403$) and outperformed the mean baseline in terms of MAE (19.6 vs. 21.7). Nevertheless, GPT-5 with search achieved substantially stronger results, indicating that while humans can do the task reasonably well, advanced reasoning models achieve markedly lower MAE. Although human forecasters might improve on this task with additional training (Tetlock et al., 2014), the fact that LLMs already outperform researchers on the performance prediction task suggests benchmark performance prediction models may provide meaningful value to practitioners when planning experiments.
6 Confidence and Retrieval Analysis
Figure 4: MAE (left axis) and coverage (right axis) as a function of confidence thresholds ($\geq$ band lower bound) for GPT-5 and Qwen3-32B, each with/without search. Higher thresholds retain fewer examples (lower coverage) and generally correspond to lower MAE.
Some experimental outcomes may be inherently difficult to predict, raising the question of whether models can reliably estimate their own prediction accuracy. To examine this, we elicit a scalar, self-reported confidence for each prediction (following Yoon et al. (2025)) and evaluate calibration by filtering to items with confidence $\geq\tau$. For each $\tau$, we report the MAE on the retained subset together with its coverage. This gives a simple view of the error–coverage trade-off. We report confidence curves for the accuracy subset, which dominates Precog.
Figure 4 shows results on the accuracy subset for GPT-5 and Qwen3-32B (with and without search). As the confidence threshold increases, coverage declines and MAE tends to decrease, indicating that higher reported confidence is, on average, associated with more accurate forecasts. Differences between “with search” and “without search” are present but modest and may vary by threshold. Overall, these curves suggest reasonable alignment between self-reported confidence and error without making strong claims about perfect calibration.
GPT5 Qwen3
Avg. search calls 3.06 0.89
Std. search calls 0.87 0.32
Unique tokens (all) 2,140 794
Unique tokens (random) 1,321
Table 6: Search-tool usage and token diversity; tokens counts measured with the GPT-4o tokenizer. (all) aggregates all formulated queries; (random) uses one random query per instance.
6.1 Retrieval Behavior
We analyze retrieval behavior via tool-call counts and query token diversity (tokenized with GPT-4o (Hurst et al., 2024)). As summarized in Table 6, GPT-5 issues more—and more variable—search calls per example than Qwen3-32B. GPT-5’s queries also show higher lexical diversity versus Qwen. These descriptive differences suggest GPT-5 tends to aggregate evidence across more—and more varied—queries, whereas Qwen’s lighter tool use may correspond to more limited gains.
7 Streaming Prediction
Using Precog with a leakage-mitigating custom retriever, GPT-5 outperforms the per-metric test-set mean baseline. Nevertheless, residual answer leakage cannot be entirely excluded—for instance, when subsequent papers copy baseline results from the original publication, or when models memorize results during pre- or post-training. To further validate our experimental setup under more realistic conditions, we adopt a streaming prediction setting, which tests models on their ability to predict performance on newly reported arXiv results immediately upon release, before they can be indexed by search engines or incorporated into training.
Setup
Figure 5: Streaming prediction error boxplots.
We collect candidates from arXiv’s official interfaces (API/RSS) for cs.CL, cs.LG, and cs.AI, and fetch PDFs via the recommended hosts, respecting robots.txt and published rate limits; we also notified arXiv staff in advance (see Appendix F). [^6] Recent lists: https://arxiv.org/list/cs.CL/recent, https://arxiv.org/list/cs.LG/recent, https://arxiv.org/list/cs.AI/recent.
A daily, low-impact pipeline then (i) identifies papers containing results that match our schema, (ii) extracts table records, (iii) builds redacted, schema-aligned descriptions (at most two per record), and (iv) issues predictions—end-to-end before third-party search engines index the papers. In this setting, we restrict to records whose metrics map to our seven metric groups (normalized to 0–100) and expand the model axis to the GPT family (beyond GPT-4/4o used on Precog). We permit GPT-5’s built-in search here, as leakage from prior literature is negligible for freshly released papers. We collected 52 papers containing 79 previously unseen experimental records and datasets between 2025-09-05 and 2025-09-23.
Results
Model MAE $\mathrm{r}$
Test set average 25.5
GPT-5 + Search 14.0 0.78
Table 7: Streaming environment prediction results on 79 samples.
Table 7 summarizes streaming performance for GPT-5 (+Search). On newly released datasets and unseen experiment–dataset combinations, the model achieves accuracy comparable to its performance on Precog (MAE =14.0). The box plot of absolute errors (Figure 5) compares all predictions with the high-confidence subset (>>50% confidence, 23 samples), showing a visibly tighter error distribution for the latter (avg. 9.3). Finally, in a self-referential test that predicted the performance of the task described in this paper (not yet on arXiv), GPT-5+Search predicted its own performance at MAE 16.0 versus the realized 14.7 from Table 3. These results indicate that forecasting performance based solely on text-only descriptions could be a practical and achievable approach in real-world applications.
8 Conclusion & Discussion
Conclusion.
We recast evaluation as forecasting and introduce Precog, a leakage-mitigating corpus of redacted, schema-aligned description–performance pairs. Search-augmented reasoning with GPT-5 produces accurate forecasts with calibrated uncertainty, and remains effective in a streaming setting where results are predicted for newly released papers prior to broad indexing. These findings indicate that anticipatory evaluation is feasible and practically useful for estimating task difficulty, prioritizing experiments, and reducing compute.
Discussion: limits of current open-source models.
Despite progress, open-source baselines lag behind GPT-5 in both accuracy and calibration. They show brittle tool use (inconsistent gains from retrieval), weaker representations of fine-grained cues in long descriptions, and higher variance across metric families (e.g., small-variance metrics like BLEU/ROUGE). Improved query planning, evidence selection, and uncertainty modeling are needed for reliable deployment.
Discussion: feasibility and utility.
[…]
Outlook: toward specialized forecasting models.
To accelerate scientific discovery, we advocate purpose-built regressors that leverage LLM reasoning. In particular, reasoning-aware regression trains a regression model on top of reasonig LLMs, while better use of retreival tool focuses on learned query planning and selective evidence use. Developing these components on top of Precog and the streaming benchmark should yield more accurate, trustworthy, and cost-efficient forecasting systems.
9 Ethics Statement
We rely exclusively on publicly available research artifacts (e.g., arXiv PDFs and published results); no human subjects or personally identifiable data are involved, and all usage follows licenses and provider terms. To reduce leakage, we redact dataset identities and block retrieval of each record’s table-source and dataset-source papers; because residual cues may persist, we also evaluate a time-gated, zero-leakage setting on genuinely unseen datasets. Focusing on arXiv means some records lack full experimental details, which can introduce label noise; to aid auditing, we release venue metadata per record and document known issues. Our strongest results use large proprietary models, raising equity and environmental concerns; we therefore report compute, provide open baselines, and encourage reproduction with smaller models. Finally, the corpus is predominantly English, limiting generality; expanding multilingual coverage is an important direction for future work.
10 Reproducibility Statement
We will release all artifacts needed for exact or statistically equivalent reproduction: the full Precog curation stack (scripts, leakage filters, redaction utilities), the forecasting pipelines (retrieval, prompting, decoding, scoring), and the evaluation scripts. Our retriever queries the arXiv API, returning at most two PDFs per query and up to four queries per instance (ReAct-style); we will also release the exact query templates and our custom arXiv wrapper. To model unseen datasets, we enforce a zero-leakage protocol by time-gating retrieval to snapshots before each dataset’s index date, and we publish per-record snapshot dates plus a script. All prompts (system/instruction/user and ReAct) are released verbatim; decoding parameters are fixed and logged, defaulting to greedy where configurable. For GPT-5, decoding settings are provider-controlled; we use defaults and record access dates. Baselines (E5-Mistral + kNN/XGBoost; Qwen3 variants) are used off-the-shelf; we provide exact checkpoints or API identifiers and access dates, and no baseline is finetuned. Upon acceptance, we will release code, data, prompts, configs, and logs under <MIT LICENSE>; non-redistributable third-party assets are referenced with exact versions and retrieval instructions.
Appendix A Details of Data Curation
This section provides data-curation details not covered in Section 4.
LLMs
For curation we used GPT-5 (OpenAI, 2025) and Qwen-32B (Yang et al., 2025). GPT-5 extracted schema-aligned descriptions from source papers; Qwen-32B handled dataset-source retrieval annotations and description rewriting. These choices were based on a pilot run and seed-result quality.
Dataset-source arXiv annotation
Given the $\mathrm{TSP}_{i}$ PDF, we prompted Qwen-32B to infer the title of the dataset-source paper, queried the arXiv API to resolve an identifier, and downloaded the PDF. Annotators then corrected errors and canonicalized dataset names.
Description extraction
We excluded appendix tables during extraction. To disambiguate records, we supplied the pre-normalization metric value as context. For efficiency, all PDFs were truncated to the first 12 pages during retrieval and extraction.
Additional filtering
All records involve GPT-4 (Achiam et al., 2023) or GPT-4o (Hurst et al., 2024), fixing the model/method axis to focus on dataset/task variation. We excluded multimodal domains, correlation-only metrics (not on a 0–100 scale), and overall benchmark aggregates (e.g., MMLU) via heuristic filters.
Appendix B Annotation Guidelines
B.1 Description Quality
This section details the human annotation protocol used to evaluate the quality of redacted, schema-aligned descriptions against their source PDFs ($\mathrm{TSP}_{i}$, $\mathrm{DSP}_{i}$). Annotators were instructed to verify that each description is anonymized, complete where possible, and strictly grounded in the provided sources.
Annotation Task
Annotators were given:
  • The extracted input description for a given experimental record.
  • The corresponding table-source and dataset-source PDFs.
The task was to check whether the description faithfully reflects information from the PDFs, is anonymized, and adheres to the fixed schema.
B.2 Evaluation Criteria
Three dimensions of quality were assessed:
Anonymization (Binary).
Descriptions must not include dataset or subset names, and must exclude performance numbers drawn from the main paper. Annotators assign either Accept (score: 1.0) or Reject (score: 0.0).
Schema coverage (1–5).
Checks whether all schema fields are filled appropriately when information is available in the PDFs, and “None” is used only when the information is truly absent.
  • Score 1: Ignores the schema; most fields missing or collapsed.
  • Score 2: Very few fields filled; many omissions; “None” often misused.
  • Score 3: About half of fields filled; some omissions or misplacements.
  • Score 4: Nearly all fields correctly covered; rare omissions or minor misplacements.
  • Score 5: Complete and consistent coverage; “None” only when truly absent; perfectly aligned with schema.
Factual grounding (1–5).
Ensures every statement in the description is directly supported by the PDFs, without speculation or outside knowledge.
  • Score 1: Mostly unsupported or invented; contradictions with PDFs.
  • Score 2: Frequent unsupported claims; weak sourcing.
  • Score 3: Generally grounded, but several unsupported claims or mild over-interpretation.
  • Score 4: Fully grounded with only minor overstatements; no contradictions.
  • Score 5: Entirely faithful; precise paraphrases of the PDFs; every fact traceable.
B.3 Human Performance Prediction
This subsection specifies the human annotation protocol for predicting task performance from redacted, schema-aligned descriptions. Annotators produce a single accuracy estimate and a brief rationale without consulting external sources, enabling an assessment of human forecasting difficulty and behavior under text-only constraints.
Annotation Task
Annotators are given:
  • A redacted description of an experimental setup (model, dataset/task, evaluation protocol, and salient configuration details).
  • A fixed-time budget per item and an interactive prompt that records their identity (first name) and assigns a shard automatically.
For each item, annotators must output:
  • Predicted accuracy (integer in 1–100).
  • Rationale (1–3 sentences) briefly justifying the estimate.
Instructions and Constraints
  • No external search. Rely solely on the provided description and personal prior knowledge; do not use web search, papers, or tools.
  • Timing. Aim to complete each prediction within 2–3 minutes.
  • Batch size. Each annotator completes 20 samples per session.
Appendix C Supplementary Experimental Results
Figure 6: Per-metric MAE (↓) by all the baselines & LLMs, grouped by evaluation metric (e.g., Accuracy, Exact (EM), Recall, F1). The red bar marks the per-metric average with its absolute MAE label; other bars show delta to that average.
Per Metric MAE for all the baselines & LLMs
We report per-metric MAE for all baselines and LLMs, along with a test-calibrated per-metric mean baseline across seven metric groups in Figure 6.
Appendix D Precog Samples
We randomly sample 5 instances from Precog for the demonstration.
Source Papers: result paper=2306.16638 Model: GPT-4 Description
Task: Binary classification of sentence pairs. Input: two short natural-language sentences presented as a pair. Label space: two classes indicating whether the second sentence logically negates the first (Yes/No; encoded as a binary output). Evaluation target: identify logical negation between sentences rather than relying on surface-level cues. Data collection: Custom corpus created by the author using designed sentence templates. For each pair, components were systematically negated (or left unmodified) to generate balanced positive and negative examples. Construction followed classical logic principles to ensure “negation” reflects logical contradiction, with emphasis on structures prone to errors (e.g., conditionals). Sentence length: approximately 5–20 words, with a roughly normal distribution. Domains: diverse everyday and professional contexts; pairs span multiple topics. Evaluation: Metric: Recall calculated for the positive class (pairs where the second sentence negates the first), comparing model predictions with gold-standard labels across the entire dataset. Difficulty characteristics: Balanced two-class classification task. Inclusion of logically complex constructions (e.g., conditional statements), rendering surface-level “not” cues unreliable. Short but variable sentence lengths (5–20 words). Designed to assess logical reasoning robustness across diverse contexts. Prompting strategy: Zero-shot classification. A fixed instruction presents the two sentences and asks whether the second negates the first, requiring a binary response (Yes/No encoded as 1/0). No in-context examples, chain-of-thought reasoning, or auxiliary decoding constraints used. Other settings: Evaluated via an API for the later-generation model in this family; each pair assessed once using the fixed prompt. No retrieval augmentation, external tools, or special truncation/windowing rules reported. Target metric: 72.8 (recall)
Source Papers: result paper=2403.00092
Model: GPT-4 Description
Task: Predict symbolic planning domain actions from natural-language procedural text. Input: a natural-language procedure together with a domain header listing available types, predicates, and the names of actions to be filled in. Output: a domain file’s action definitions; each action must include typed parameters, a conjunction of preconditions, and a conjunction of effects, expressed in a formal planning-language syntax. Constraint: models must not add, remove, or rename actions. Evaluation target: reproduce the gold action definitions so that the predicted domain file matches the reference. Data collection: Source texts are open-domain how-to guides from the web covering varied everyday and survival-style topics. For each procedure, trained experts manually created paired symbolic planning representations: a domain file (types, predicates, actions) and associated problem files (entities, initial and goal states). Representations are planner-ready and not tied to any specific simulator. Evaluation: Metric: action-level accuracy under intrinsic evaluation. Protocol: compare predicted actions with the reference actions in the gold domain file using a semantic-equivalence check that disregards variable renaming and conjunct order; parameters must match by type, and predicates must align in arity and argument mapping. Subset: intrinsic—only direct semantic alignment of predicted action definitions is assessed; no plan execution/solvability is considered. Difficulty characteristics: Open-domain content with diverse topics. Requires extracting relevant information, inferring implicit entity states, and precisely translating into a low-resource, domain-specific formal language. Predicting preconditions is generally more challenging than predicting effects. Input texts can be long (from concise summaries to multi-paragraph procedures), creating context-length and information-extraction challenges. Prompting strategy: Zero-shot direct instruction without step-by-step reasoning scaffolding. The prompt provides the domain header (types, predicates, action names) and the relevant procedure text, and asks the model to complete each action’s parameters, preconditions, and effects in the required formal syntax while preserving the given action set. Other settings: Evaluated via an API for GPT-4; the input budget was capped at roughly ten thousand tokens for these runs. Target metric: 15.9 (accuracy)
Source Papers: result paper=2310.11191, dataset paper=2104.05767
Model: GPT-4 Description
Task: Paragraph-to-document–level text simplification in the medical domain. Input: one or more paragraphs from technical abstracts of clinical evidence reviews. Output: a plain-language, open-ended rewrite that retains key findings. Answer space: unrestricted natural language (no predefined options). Evaluation target: alignment of the generated simplification with the reference plain-language version in content and lexical overlap. Data collection: Source texts: technical abstracts from evidence syntheses on clinical topics; targets: plain-language summaries authored by domain experts for non-expert audiences. English-only; sourced from a large online repository of medical systematic reviews. Abstracts and corresponding lay summaries were paired via heuristic section matching: results-focused abstract sections were retained; for lay summaries, sections discussing studies/results (or the first paragraph containing terms like “study/trial” for unstructured summaries) were selected. Pairs with large length discrepancies were removed; examples were filtered to fit standard transformer context budgets. Note: lay summaries are not direct simplifications of abstracts (derived from full reviews), yielding partial content overlap rather than sentence-level parallelism. Evaluation: Metric: ROUGE-LSum (F1, longest common subsequence at the summary level) between model output and the single reference lay summary. Difficulty characteristics: Domain shift and style gap: targets are stylistically simpler but only marginally easier on readability metrics; success requires controlled paraphrasing and style adaptation beyond sentence/word shortening. Technical jargon and statistical elements (e.g., measurements, confidence intervals) increase lexical/semantic complexity. Non-parallelity: targets may include information absent from abstracts (e.g., currency statements), creating source–target mismatches and risk of hallucination. Long, multi-paragraph inputs necessitate document-level content selection and abstraction. Prompting strategy: Zero-shot instruction prompting: a system instruction specifies a simplification assistant role; a user instruction asks to simplify the provided text. No examples or chain-of-thought demonstrations are provided. Other settings: No retrieval augmentation, external tools, or special decoding constraints reported for this record. Target metric: 33.9 (rouge)
Source Papers: result paper=2310.20105
Model: GPT-4 Description
Task: Single-label, multi-class text classification of student help requests from introductory programming courses. Input: a natural-language issue description, optionally including a code snippet and an error message; the prompt also provides a numeric indicator of alignment between the message and assignment instructions. Label space: six mutually exclusive categories—three debugging subcategories (error message present; expected outcome specified; both present), plus implementation, understanding, and a nothing/insufficient-content class. Target: assign a single category to each request. Data collection: Source: requests submitted via a course help assistant using a semi-structured form. Captured contents: issue description, optional code and error message (programming language collected by the form but not used in the prompt). Annotation: two human coders independently labeled requests; disagreements were resolved by adopting the more experienced coder’s label; inter-rater agreement reported as substantial. Domain/language/licensing: programming-education context; language unspecified; no licensing details provided. Evaluation: Metric: F1 score for the specific category (per-class F1) on the test set, comparing model predictions with human reference labels. Subset: Implementation category only—requests seeking guidance on implementing code for a specific assignment problem (often include code and/or explicit references to instructions). Difficulty characteristics: Six-way classification with nuanced distinctions based on subtle cues (e.g., presence/absence of code, explicit error messages vs. desired outcomes, references to instructions). Inputs may include code and error traces; no prompt-length constraints were reported to affect evaluation. Prompting strategy: Zero-shot classification (no in-context examples, no chain-of-thogught). System instruction outlines the full coding scheme and decision criteria (adapted from human-coder guidelines) and requires output of a single category label without explanation. User message includes: the student’s issue description, any provided code, any provided error message, and a numeric “issue–instructions equivalence” percentage indicating overlap with assignment instructions (% only used as context). Other settings: Inference via API, one request at a time. temperature 0.0, top_p 1, max_tokens 10, frequency_penalty 0, presence_penalty 0. Target metric: 85 (F1)
Source Papers: result paper=2406.01238, dataset paper=1803.06643
Model: GPT-4 Description
Task: Open-domain natural-language question answering over knowledge graphs. Input: a single English question; the system may decompose it into sub-questions and navigate a knowledge graph. Output: one or more target entities/values from the graph; the final prediction is the top candidate. Evaluation target: whether the top predicted answer matches any reference answer. Data collection: Built by generating complex queries from a simpler QA corpus that pairs questions with graph queries over a large collaborative knowledge graph. Complexity introduced via function composition, conjunctions, superlatives, and comparatives. Machine-generated templates derived from graph predicates were paraphrased by crowd workers into natural language; gold answers obtained by executing graph queries. Domain: broad, open-domain factual knowledge; Language: English; Provenance: curated knowledge graph + crowdsourced paraphrases. Evaluation: Metric: answer accuracy (Hits@1), i.e., whether the top prediction matches any gold answer string/entity. Difficulty characteristics: Multi-step reasoning with compositions and set operations (e.g., conjunctions), plus superlative/comparative constraints. Paraphrase variability: human rewrites introduce synonymy, reordering, omissions, and additions. Open-domain coverage and entity-centric reasoning increase ambiguity, requiring precise entity typing and relation selection. Often requires multi-hop traversal over the graph to reach the final answer. Prompting strategy: Strategic, iterative LLM-driven planning over a knowledge graph (EffiQA). The high-capacity LLM performs: 1) global planning to decompose the question and generate exploration instructions (with simulated answers to guide search); 2) delegated efficient KG exploration via breadth-first expansion with semantic pruning using fine-grained entity typing and similarity to simulated answers; 3) self-reflection to review partial results, revise the plan, and iterate until sufficient evidence is gathered. Other settings: LLM temperature set to 0 for deterministic outputs. Exploration uses breadth-first search with thresholds to control branching; relations/entities not matching planned types/qualifiers are pruned. A plug-in encoder-based classifier (fine-grained entity typing) supports constrained semantic matching to maintain recall while reducing search space. Stopping criterion: output an answer when results are satisfactory; otherwise replan and iterate (self-reflection loop). Target metric: 69.5 (accuracy)
Appendix E Prompts
E.1 Prompt for prediction using search
Prompt
Given an experimental record (task, model, and setup), search for supporting evidence and estimate the LLM’s performance for the specified setup. Inputs
Target model: model evaluated on. Metric family: metric family used to evaluate the experimental record. Redacted description: description of the experimental record (e.g., task, dataset collection, evaluation metric, experimental setup). Search Process (REQUIRED)
Use the search tool up to four times to gather evidence. Keep queries concise; the retrieval module cannot handle complex queries. Focus queries on the Target model evaluated on related datasets/tasks (e.g., GPT-4o clinical summarization rouge). Submit at most one query per turn; after each query, the tool returns matching paper PDFs. Plan what to search before calling the tool. If retrieved evidence is irrelevant, do not reference it for the prediction. Confidence Assessment
After producing a performance prediction, critically review your reasoning and the evidence considered. Classify confidence into one of: Almost no chance, Highly unlikely, Chances are slight, Unlikely, Less than even, Better than even, Likely, Very good chance, Highly likely, Almost certain. Output Format
[…]
E.2 Prompt for prediction without using search
Prompt
Given an experimental record (task, model, and setup), estimate the LLM’s performance for the specified setup. Inputs
Target model: model evaluated on. Metric family: metric family used to evaluate the experimental record. Redacted description: description of the experimental record (e.g., task, dataset collection, evaluation metric, experimental setup). Confidence Assessment
After producing a performance prediction, critically review your reasoning and the evidence considered. Classify confidence into one of: Almost no chance, Highly unlikely, Chances are slight, Unlikely, Less than even, Better than even, Likely, Very good chance, Highly likely, Almost certain. Output Format
[…]
Appendix F Compliance with arXiv Access Policies
We accessed arXiv content in accordance with arXiv’s published guidance for automated use, including the robots.txt [^7] https://arxiv.org/robots.txt and the “Robots Beware” help page. [^8] https://info.arxiv.org/help/robots.html Our goal was to minimize load on arXiv’s infrastructure while ensuring traceable, policy-compliant retrieval.
Scope of access.
Agents fetched only URLs permitted by robots.txt and avoided disallowed paths. We did not perform indiscriminate crawling or attempt corpus-scale downloads via the human-facing site.
Preferred interfaces.
Wherever possible, we used arXiv’s programmatic interfaces (e.g., API/OAI-PMH/RSS) instead of scraping HTML pages, consistent with the help page’s guidance to favor machine interfaces for automated access.22footnotemark: 2
Robots rules and rate limiting.
Our clients honored robots.txt directives (allow/disallow and crawl-delay) and enforced conservative global throttling with backoff on errors. We limited parallelism and maintained a single logical connection per agent run.
Identification and advance notice.
All requests included a descriptive User-Agent string and a monitored contact email, as recommended.22footnotemark: 2 Prior to large-scale experiments, we emailed the arXiv team responsible for automated access to describe our usage pattern and confirm that no special coordination was required.
Low-impact retrieval policy.
To further reduce load, each retrieval-capable agent capped (i) the number of queries it could issue and (ii) the number of documents returned per query. We cached responses, deduplicated URLs, and normalized links to canonical arXiv endpoints to avoid redundant traffic.
Monitoring and fail-safe behavior.
We continuously monitored HTTP status codes and immediately slowed or halted traffic upon receiving 429/403/5xx responses. Our clients also rechecked robots.txt at regular intervals to respect policy changes; any change that tightened restrictions triggered an automatic stop.
Summary.
In aggregate, our practices align with arXiv’s automated access rules: we honored robots.txt, preferred official machine interfaces, rate-limited conservatively, identified ourselves, notified arXiv in advance, and implemented additional safeguards to minimize load.
Appendix G LLM Usage Disclosure
We disclose all uses of large language models (LLMs) in the preparation of this work.
Scope of use
LLMs were used only for (i) data curation utilities and (ii) baseline models in our experiments, and (iii) light writing assistance. LLMs were not used for research ideation, experimental design, claim formation, or result interpretation.
Data curation
LLMs assisted with mechanical curation tasks such as dataset source annotation, extraction, and rewriting. All curation outputs were reviewed by the authors; any labels or annotations used for evaluation were not generated solely by LLMs.
Baselines
As reported in the main text (e.g., table 3), we evaluate several off-the-shelf LLMs strictly as baselines. These models were used via their standard inference APIs with fixed prompts and hyperparameters (temperature, top-$p$, max tokens) specified in the paper/appendix. No additional finetuning of these baseline LLMs was performed.
Writing assistance
LLMs were used for copy-editing only: grammar checks, minor rephrasing of awkward sentences, and generating LaTeX boilerplate (e.g., table skeletons). All scientific content, analysis, and conclusions are authored by the human authors; every LLM-suggested edit was manually reviewed and, when necessary, revised.
Non-use
LLMs were not used for ideation or synthesis of contributions. They did not select experiments, decide hypotheses, or write technical sections beyond copy-editing as described above.
Privacy and safety
No private or personally identifiable information was provided to LLMs. Inputs were limited to public research materials and our own synthetic/derived artifacts. API usage complied with the respective providers’ terms of service.
Author responsibility
The authors take full responsibility for the paper’s content. LLM assistance was limited to the bounded uses above and did not alter the scientific claims or their justification.